3. Show that the function f(x, y) = V26 – 2x2 - y2 is continuous at point (2, -3).
Answer #1

ya f(x1=(26-2x²y² P=(2, -3) replace x=2+rcond y=-3tr sind fir,dl= √26-2(2+0000²-(-3+2Sinaj ? = $26-2 (4+84220 +4688)-(9 trsin